A rectangular plate has a length of 21.3
1
Expert's answer
2021-10-11T16:36:27-0400

"a=(21.3\u00b10.2)~cm,"

"b=(9.8\u00b10.1)~cm,"

"A=ab=21.3\cdot 9.8±21.3\cdot 0.1±9.8\cdot 0.2+0.2\cdot 0.1=(208.76±4.09)~cm^2."